OpenStudy (anonymous):

Algebra 1 help? X + 1 divided by 3 = 6

OpenStudy (marissalovescats):

You multiply both sides by 3 to cancel the 3 which leaves you with x+1=6(3) which is x+1=18 and then you subtract 1 from each side to get x=17

OpenStudy (anonymous):

When you multiply both sides by 3, does that not leave you with 3x + 3 = 18, subtract 3 from both sides and divide by 3 to get x = 5

OpenStudy (anonymous):

3(x + 1) = 6(3)

OpenStudy (anonymous):

I may be wrong which is why I was checking

OpenStudy (marissalovescats):

No